OpenStudy (anonymous):

Complete the table and choose the solution for each given value of x in questions 14–18. EQUATION IN REPLY DEAL WITH GRAPHS IF WE GET THERE 14. When x = –4 (1 point) 15. When x = –2 (1 point) 16. When x = 0 (1 point) 17. When x = 2 (1 point) 18. When x = 4 (1 point) 19. Select the graph that matches the function. (2 points)

OpenStudy (anonymous):

\[f(x)=\frac{1}{2}|x|-4\]?

OpenStudy (anonymous):

Yes that seems to be the equation without the f (x)

OpenStudy (anonymous):

\[f(-4)=\frac{1}{2}|-4|-4=\frac{1}{2}\times 4-4=2-4=-2\] the rest are similar

OpenStudy (anonymous):

you will notice that \(x=-4\) and \(x=4\) give the same answers because \(|-x|=|x|\)

OpenStudy (anonymous):

correct

OpenStudy (anonymous):

imagine graph looks like this http://www.wolframalpha.com/input/?i=1%2F2 |x|-4

OpenStudy (anonymous):

cut and paste, absolute value reeks havoc on links

OpenStudy (anonymous):

yes I have noticed that

OpenStudy (anonymous):

so number 14 would be https://www.connexus.com/content/media/149657-2162007-30814-PM-827475187.gif